# Efficient and Accurate Optimal Transport with Mirror Descent and Conjugate Gradients

### v3

@@ 4.2 A mirror descent method for optimal transport: MDOT
$t\leftarrow 1$, $\textrm{done}\leftarrow\textrm{False},\gamma\leftarrow\gamma_{\mathrm{i}}\wedge\gamma_{\mathrm{f}}$
2:while not done do
3:     $\textrm{done}\leftarrow\gamma==\gamma_{\mathrm{f}}$
4:     $\varepsilon_{\mathrm{d}}\leftarrow H_{\min}({\bm{r}},{\bm{c}})/\gamma^{p}$
5:     $(\tilde{{\bm{r}}},\tilde{{\bm{c}}})\leftarrow(1-\frac{\varepsilon_{\mathrm{d}}}{4})\cdot({\bm{r}},{\bm{c}})+\frac{\varepsilon_{\mathrm{d}}}{4n}\cdot\bm{1}_{2n}$
6:     if $t==1$ then ${\bm{z}}^{(t)}\leftarrow(\log\tilde{{\bm{r}}},\log\tilde{{\bm{c}}})$
7:     ${\bm{z}}^{(t)}\leftarrow\textrm{BregmanProject}({\bm{z}}^{(t)},\gamma,C,\tilde{{\bm{r}}},\tilde{{\bm{c}}},\varepsilon_{\mathrm{d}}/2)$
8:     $\gamma\leftarrow q\gamma\wedge\gamma_{\mathrm{f}}$
9:     ${\bm{z}}^{(t+1)}\leftarrow\textrm{WarmStart}({\bm{z}}^{(t)};\cdots)$
@@
12:$({\bm{u}},{\bm{v}})\leftarrow{\bm{z}}^{(t-1)}$, $P\leftarrow\exp\{{\bm{u}}\bm{1}_{n}^{\top}+\bm{1}_{n}{\bm{v}}^{\top}-\gamma_{\mathrm{f}}C\}$
@@
Here we introduce the generic MDOT framework shown in Alg. 1 which will be exemplified throughout this section. Note that a routine in L9 is defined using $\cdots$” as a placeholder for additional parameters that may be required by specific implementations. MDOT accepts as input problem parameters $C,{\bm{r}},{\bm{c}}$, as well as user-provided positive scalars $\gamma_{\mathrm{i}},\gamma_{\mathrm{f}},p,q$. Here, $\gamma_{\mathrm{i}}$ will be the initial step size $\Delta_{\gamma}^{(0)}=\gamma^{(1)}$ of the (inexact) mirror descent procedure mimicking (6), while $\gamma_{\mathrm{f}}$ will be the final (inverse) temperature. As seen in L8, near the end of each outer loop iteration, MDOT decays the temperature under a schedule $\gamma^{(t+1)}=q\gamma^{(t)}$ until $\gamma^{(t)}=\gamma_{\mathrm{f}}$ similarly to Schmitzer (2019); Feydy (2020). Observe that due to L1 and L3, if called with $\gamma_{\mathrm{i}}\geq\gamma_{\mathrm{f}}$, MDOT terminates after a single iteration of the while loop.
At each value of the temperature, in L4 we first select a tolerance $\varepsilon_{\mathrm{d}}=H_{\min}({\bm{r}},{\bm{c}})/\gamma^{p}$ for dual objective gradient norm $\left\lVert\nabla g\right\rVert_{1}$, where ${p\geq 1}$. This is in contrast to the more standard usage $\varepsilon_{\mathrm{d}}\propto(\log n)/\gamma$ in prior work. In Appx. C, we show the advantage of using $H_{\min}({\bm{r}},{\bm{c}})$ over $\log n$ empirically (for $p=1$), demonstrating substantial speedups of order $\log n/H_{\min}({\bm{r}},{\bm{c}})$ for Sinkhorn iteration. The case of $p>1$ is discussed in detail in Sec. 4.2.2. In L5 of Alg. 1, we “smooth” the target marginals ${\bm{r}}$ and ${\bm{c}}$ by mixing in uniform distributions, similarly to prior work by Dvurechensky et al. (2018) and Lin et al. (2019). This step helps provide convergence guarantees for certain choices of minimization algorithms for problem (2.1) used in L7. Since the mixing weight used in L5 is proportional to $\gamma^{-p}$, i.e., it gradually decreases with the temperature, this scheme smoothes marginals more aggressively in earlier iterations of MDOT. In Appx. D, we empirically show the performance benefits of this variable smoothing.
In L6, the dual variables are initialized as ${{\bm{u}}=\log\tilde{{\bm{r}}}}$ and ${{\bm{v}}=\log\tilde{{\bm{c}}}}$, which corresponds to initializing ${\tilde{P}^{(0)}=\tilde{{\bm{r}}}\tilde{{\bm{c}}}^{\top}\in{\mathcal{U}}(\tilde{{\bm{r}}},\tilde{{\bm{c}}})}$, i.e., the maximum entropy plan in ${\mathcal{U}}(\tilde{{\bm{r}}},\tilde{{\bm{c}}})$, which is rank-1 as required by Prop. 4.1. Next in L7, given the initial guess for the dual variables ${\bm{z}}=({\bm{u}},{\bm{v}})$ the objective in (2.1) is minimized under the smoothed marginals $\tilde{{\bm{r}}},\tilde{{\bm{c}}}$ until $\left\lVert{\bm{r}}(P)-\tilde{{\bm{r}}}\right\rVert_{1}+\left\lVert{\bm{c}}(P)-\tilde{{\bm{c}}}\right\rVert_{1}\leq\varepsilon_{\mathrm{d}}/2$. This step is called BregmanProject as it carries out (inexact) Bregman projection of the plan $P({\bm{u}}^{(t)},{\bm{v}}^{(t)};\gamma)$ onto ${\mathcal{U}}(\tilde{{\bm{r}}},\tilde{{\bm{c}}})$ as in (5). The algorithm to be used for minimizing the dual objective is left unspecified here for generality; for example, Sinkhorn iteration can be used. In 4.3, we introduce a new algorithm (PNCG) for this purpose. L8 sets $\gamma^{(t+1)}=q\gamma^{(t)}$, which corresponds to a default mirror descent step size $\Delta_{\gamma}=(q-1)\gamma$. In L9 we use the current near-optimal dual variable ${\bm{z}}^{(t)}$ (and perhaps additional ${\bm{z}}^{(t-k)}$’s for $k\geq 1$) to produce a warm-starting guess for the optimal dual variables at the next temperature value. Finally in L13, after an approximation $P$ of $P^{*}(\gamma_{\mathrm{f}})$ has been obtained, $P$ is rounded onto ${\mathcal{U}}({\bm{r}},{\bm{c}})$ via Alg. 2 of Altschuler et al. (2017); see Alg. 3 in Appx. A.
@@
If $\gamma_{\mathrm{f}}=5H_{\min}\big{(}{\bm{r}},{\bm{c}}\big{)}/2\varepsilon$, the output $P\in{\mathcal{U}}({\bm{r}},{\bm{c}})$ of Alg. 1 satisfies $\langle P-P^{*},C\rangle\leq\varepsilon+\widetilde{O}(\varepsilon^{2})$.
Note that the above guarantee can be obtained considering only the very last iteration of MDOT. That is, the specific assignments in L4-5 in prior iterations, choices $p,q$ and implementations of L7 and L9 play a significant role on the overall runtime of the algorithm (as we shall see), but not the guaranteed precision of the final solution returned as per Remark 4.2 above.
In the remainder of this section, we will:
  1. 1.
    Develop a warm-starting routine and compare it to a prior approach in temperature annealing.
  2. 2.
    Investigate the implications of various choices for the $p$ and $q$ parameters via ablation experiments.
  3. 3.
    Show the potential advantage of choosing $p>1$ theoretically (Prop. 4.4).
  4. 4.
    Investigate a new BregmanProject routine based on non-linear CG for problem (2.1) in L7 of MDOT.
4.2.1 Warm-starting Bregman Projections
@@
As we cannot compute $\partial{\bm{z}}^{*}/\partial\gamma$ analytically, we use a numerical approximation (backward finite differencing):
$\frac{\partial{{\bm{z}}^{*}}}{\partial\gamma}(\gamma^{(t)})\approx\frac{{\bm{z}}^{*}(\gamma^{(t)})-{\bm{z}}^{*}(\gamma^{(t-1)})}{\gamma^{(t)}-\gamma^{(t-1)}}.$ (10)
Keeping the first two terms in (9) and rearranging:
@@
Note that in practice, optimal ${\bm{z}}^{*}$ are unknown; only approximate minimizers are available. Since the Taylor approximation and backward finite differencing in (10-11) already incur some amount of error, care should be taken not to over-optimize (2.1) in an effort to better approximate ${\bm{z}}^{*}$ in intermediate steps $t$.
@@
4.2.2 Bregman Projection Stopping Criteria
Our assignment $\varepsilon_{\mathrm{d}}\propto\gamma^{-p}$ for some $p\geq 1$ in L4 of MDOT departs from the conventional wisdom of choosing $\varepsilon_{\mathrm{d}}\propto\gamma^{-1}$ (Altschuler et al., 2017; Dvurechensky et al., 2018; Lin et al., 2019). Here, we aim to provide justification for this departure. Consider first the fixed-temperature problem (2.1) for simplicity. building on the results of Cominetti & Martín (1994), Weed (2018) showed in his Prop. 4 and Thm. 5 that there is both a uniform bound $\langle P^{*}(\gamma)-P^{*},C\rangle\leq\log n/\gamma$ (slow rate), and a fast asymptotic rate $O(\exp(-\gamma K))$ which takes over for large enough $\gamma$, where the constant $K>0$ is problem-dependent. Taking these as a starting point, the following remark generalizes the third statement of Prop. 4.1.
@@ 4.2.2 Bregman Projection Stopping Criteria
Figure 2: Ablation of stopping criterion parameter $p$ (Sec. 4.2.2) for the SK algorithm (left) and MDOT with Sinkhorn and PNCG as Bregman projectors (right). The SK algorithm (left) is called by running MDOT (Alg. 1) with $\gamma_{\mathrm{i}}=\gamma_{\mathrm{f}}$, where higher precision is achieved by increasing $\gamma_{\mathrm{f}}$. Results show the median over 36 random problems from the upsampled MNIST dataset ($n=4096$) with the $L_{1}$ cost.
This result is consistent with the empirical findings of Jambulapati et al. (2019), who noted “The [tuned] Sinkhorn algorithm converged at rates much faster than the predicted $\varepsilon^{-2}$ rate on all experiments, outperforming all other methods, which we believe merits further investigation.” We believe Prop. 4.4 sheds some light on this phenomenon, and further present an ablation of $p$ in Fig. 2 for SK and MDOT algorithms.
For the SK algorithm, Fig. 2 (left) verifies the insight derived from Prop. 4.4. The choice $p=1$ is better at low precision, but the trend gradually shifts in favor of higher $p$ with (sufficiently) higher $\gamma_{\mathrm{f}}$. That is, for sufficiently low temperature $\gamma^{-1}$, it is advantageous to reduce the gradient norm error tolerance, from $H_{\min}/\gamma$ to $H_{\min}/\gamma^{p}$ for $p>1$. In contrast, MDOT is more robust to the $p$ parameter in the high precision regime (right). Moreover, the use of PNCG projections (Alg. 2) for Bregman projections in MDOT (L7 of Alg. 1) provides a speedup of 23×2-3$\times$2 - 3 × over Sinkhorn projections. PNCG is introduced and discussed next.
4.3 Preconditioned Non-linear Conjugate Gradients for Bregman Projections
Algorithm 2 PNCGProject$({\bm{z}},\gamma,C,{\bm{r}},{\bm{c}},\varepsilon_{\mathrm{d}})$

### v4

@@ 4.2 A mirror descent method for optimal transport: MDOT
Using insights derived in the previous section, we construct the MDOT method shown in Alg. 1. As illustrated in Fig. 1, MDOT minimizes a sequence of EOT dual objectives $g(\gamma^{(t)})$. At each step, MDOT uses prior approximate minimizers ${\bm{u}}^{(t-k)},{\bm{v}}^{(t-k)}$ of $g(\gamma^{(t-k)})$ for $k\in[0,K]$ to extrapolate better initial guesses (warm-starting) for the next problem at $t+1$ (see Fig. 1, where $K=1$ is depicted).
Figure 1: In each iteration of Alg. 1, MDOT approximately minimizes $g({\bm{u}},{\bm{v}};\gamma)$ as in (8) to compute duals $({\bm{u}}^{(t)},{\bm{v}}^{(t)})$ (i.e., “MDOT iterates”), starting initially with $\gamma_{\mathrm{i}}$ and increasing $\gamma$ by a factor of ${q>1}$ per iteration until a final value $\gamma_{\mathrm{f}}$. With increasing $\gamma$ (weaker entropic regularization), the convex dual objectives defined over $\mathbb{R}^{2n}$ become increasingly ill-conditioned (illustrated here in 1D by parabolas with increasingly high curvature) and therefore harder to solve. At each step, MDOT uses approximate solutions of prior problems to produce an initial guess for the next problem (left), and generates increasingly accurate warm-start guesses to mitigate the difficulty posed by ill-conditioning (right).
The MDOT loop is summarized with the following steps:
  • L3: Choose stopping criterion $\varepsilon_{\mathrm{d}}$ for $\left\lVert\nabla g(\gamma,{\bm{r}},{\bm{c}})\right\rVert_{1}$.
  • L4: Obtain smoothed marginals $\tilde{{\bm{r}}},\tilde{{\bm{c}}}$ to avoid zero or infinitesimal entries.
  • L5: Rank-1 initial guess with positive $\tilde{{\bm{r}}},\tilde{{\bm{c}}}$.
  • L6: Solve (8) given initial ${\bm{u}}^{\prime},{\bm{v}}^{\prime}$, i.e., minimize $g(\gamma,\tilde{{\bm{r}}},\tilde{{\bm{c}}})$ until ${\left\lVert\nabla g(\gamma,\tilde{{\bm{r}}},\tilde{{\bm{c}}})\right\rVert_{1}\leq\varepsilon_{\mathrm{d}}/2}$. By the triangle inequality, this implies ${\left\lVert\nabla g(\gamma,{\bm{r}},{\bm{c}})\right\rVert_{1}\leq\varepsilon_{\mathrm{d}}}$.
  • L7: Exit if $\gamma$ reached user input $\gamma_{\mathrm{f}}$, otherwise continue.
  • L8: Increase $\gamma$ for the next iteration.
  • L9: Using previous ${\bm{u}}^{(t-k)},{\bm{v}}^{(t-k)}$ for $k\geq 0$, extrapolate a warm-start guess for the new KL projection (or, EOT dual) problem of minimizing $g(\gamma^{(t+1)},{\bm{r}},{\bm{c}})$.
In L13, approximation of $P^{*}(\gamma_{\mathrm{f}})$ is rounded via Altschuler et al. (2017) onto ${\mathcal{U}}({\bm{r}},{\bm{c}})$; see Alg. 3 in Appx. A.
@@
$t\leftarrow 1$, $\gamma\leftarrow\min(\gamma_{\mathrm{i}},\gamma_{\mathrm{f}})$
2:while True do
3:     $\varepsilon_{\mathrm{d}}\leftarrow H_{\min}({\bm{r}},{\bm{c}})/\gamma^{p}$
4:     $(\tilde{{\bm{r}}},\tilde{{\bm{c}}})\leftarrow(1-\frac{\varepsilon_{\mathrm{d}}}{4})\cdot({\bm{r}},{\bm{c}})+\frac{\varepsilon_{\mathrm{d}}}{4n}\cdot\bm{1}_{2n}$
5:     if $t==1$ then ${\bm{u}}^{\prime},{\bm{v}}^{\prime}\leftarrow\log\tilde{{\bm{r}}},\log\tilde{{\bm{c}}}$
6:     ${\bm{u}}^{(t)},{\bm{v}}^{(t)}~{}{\leftarrow}~{}\textrm{Given initialization ${\bm{u}}^{\prime},{\bm{v}}^{\prime}$}\linebreak\textrm{minimize }g(\gamma,\tilde{{\bm{r}}},\tilde{{\bm{c}}})\textrm{ until }\left\lVert\nabla g\right\rVert_{1}\leq{\varepsilon_{\mathrm{d}}/2}$
[…]
8:     $\Delta_{\gamma}\leftarrow(q-1)\gamma,\quad\gamma\leftarrow\min(\gamma+\Delta_{\gamma},\gamma_{\mathrm{f}})$
9:     ${\bm{u}}^{\prime},{\bm{v}}^{\prime}\leftarrow\textrm{WarmStart}({\bm{u}}^{(t)},{\bm{v}}^{(t)};\cdots)$
@@
12:$P\leftarrow\exp\{{\bm{u}}^{(t)}\bm{1}_{n}^{\top}+\bm{1}_{n}{{\bm{v}}^{(t)}}^{\top}-\gamma_{\mathrm{f}}C\}$
@@
Using (i) the guarantee that $\left\lVert\nabla g(\gamma_{\mathrm{f}},{\bm{r}},{\bm{c}})\right\rVert_{1}\leq H_{\min}({\bm{r}},{\bm{c}})/\gamma_{\mathrm{f}}$, (ii) the third statement of Prop. 4.2, and (iii) Lemma 7 of Altschuler et al. (2017), we obtain a guarantee on the quality of the solution.
@@
{mdframed}
[ skipabove=5pt, skipbelow=0pt, innertopmargin=2pt, innerbottommargin=2pt] If $\gamma_{\mathrm{f}}\geq 5H_{\min}\big{(}{\bm{r}},{\bm{c}}\big{)}/2\varepsilon$, the output $P\in{\mathcal{U}}({\bm{r}},{\bm{c}})$ of Alg. 1 satisfies $\langle P-P^{*},C\rangle\leq\varepsilon+\widetilde{O}(\varepsilon^{2})$.
The proof follows from a special case of Prop. 4.5, which is deferred to Sec. 4.2.2. Next, we discuss key MDOT steps listed above, how they differ from existing work, and trade-offs of input parameters.
Stopping criterion (L3) and input $p$. Several prior methods such as IPOT (Xie et al., 2020), Alg. 3.5 of Feydy (2020) and MSK (Ballu & Berthet, 2023) that use MD ideas do not enforce any constraints on $\left\lVert\nabla g\right\rVert_{1}$. Rather, they only take a few Sinkhorn steps after $\gamma$ is increased. Feydy (2020) takes increasingly bigger updates $\Delta_{\gamma}=(q-1)\gamma$ as we do in L8, such that when $\gamma$ is large, a few Sinkhorn steps are insufficient to prevent divergence from the optimal curve ${\bm{u}}^{*}(\gamma),{\bm{v}}^{*}(\gamma)$. This effectively caps their precision as we show in Sec. 5. IPOT takes fixed steps (typically $\Delta_{\gamma}=1$), which helps trace the curve ${\bm{u}}^{*}(\gamma),{\bm{v}}^{*}(\gamma)$ closely, but reduces the regularization weight $\gamma^{-1}$ more slowly; see Appx. G for further discussion and empirical data. On the other hand, MSK takes increasingly smaller $\Delta_{\gamma}^{(t)}=O(1/\sqrt{t})$, such that its convergence rate suffers (shown in Sec. 5). All three rely on Sinkhorn updates, while MDOT can benefit from any (potentially better) convex optimization algorithms for minimizing $g$. The practical approach of Yang & Toh (2022) also uses a fixed $\Delta_{\gamma}$ and Sinkhorn iteration. As the stopping criterion, they measure $D_{\mathrm{KL}}(\mathrm{Round}(P,{\bm{r}},{\bm{c}})|P)$ after each Sinkhorn step to satisfy an inexactness condition for the KL projection, which facilitates their theoretical analysis. This incurs some $O(n^{2})$ overhead per inner loop iteration. In contrast, we only check the $L_{1}$ norm of the dual gradient as in L6 (which can be evaluated in $O(n)$ time during optimization), since by Proposition 4.2 maintaining the structure $P({\bm{u}},{\bm{v}};\gamma)$ throughout ensures that $P^{*}(\gamma_{\mathrm{f}})$ can be recovered theoretically, regardless of how inexact previous KL projections were.
Another line of work runs (some version of) a single iteration of MDOT (Altschuler et al., 2017; Dvurechensky et al., 2018; Lin et al., 2019). In these works, $\varepsilon_{\mathrm{d}}\propto(\log n)/\gamma$, while we adapt it according to the entropy of ${\bm{r}},{\bm{c}}$ using the third statement of Prop. 4.2. In Appx. C, the use of $H_{\min}({\bm{r}},{\bm{c}})$ instead of $\log n$ is shown to yield speedups of order $\log n/H_{\min}({\bm{r}},{\bm{c}})$. Our use of $p\geq 1$ is also novel and controls how closely the curve ${\bm{u}}^{*}(\gamma),{\bm{v}}^{*}(\gamma)$ is traced. We study theoretical benefits and practical trade-offs in Sec. 4.2.2.
Marginal smoothing (L4). Our “smoothing” of the target marginals ${\bm{r}}$ and ${\bm{c}}$ (mixing in the uniform distribution) follows prior work by Dvurechensky et al. (2018) and Lin et al. (2019). This step helps provide convergence guarantees for certain choices of minimization algorithms for the dual $g$ in L6. Since the mixing weight used is proportional to $\gamma^{-p}$, it gradually decreases with the temperature in our case. This scheme smoothes marginals more aggressively in earlier iterations of MDOT when $\gamma$ is small. In Appx. D, we empirically show the performance benefits of this variable smoothing.
EOT dual minimization (L6). The algorithm for minimizing $g(\gamma,{\bm{r}},{\bm{c}})$ is left unspecified here for generality; for example, Sinkhorn iteration can be used. In Sec. 4.3, we introduce a new algorithm (PNCG).
Temperature decay and input $q$ (L8).
L8 effectively sets $\gamma^{(t+1)}=q\gamma^{(t)}$, which follows prior work (Schmitzer, 2019; Feydy, 2020). With larger MD steps (higher $q$), the extrapolation of duals for warm-starting is over a longer range, which degrades quality and poses a trade-off, investigated in Sec. 4.2.1.
Warm-starting duals (L9). Our explicit warm-starting approach in Sec. 4.2.1 is novel to our knowledge and was left unspecified in Alg. 1 for generality. We show that $\varepsilon$-scaling as implemented by Schmitzer (2019) and Feydy (2020) amount to an implicit warm-starting, which is shown to be less effective than ours.
4.2.1 Warm-starting KL Projections
@@
As we cannot compute $\partial{\bm{z}}^{*}/\partial\gamma$ analytically, we use a numerical approximation (backward finite differencing) $\partial{{\bm{z}}^{*}}(\gamma^{(t)})/\partial\gamma\approx\big{(}{\bm{z}}^{*}(\gamma^{(t)})-{\bm{z}}^{*}(\gamma^{(t-1)})\big{)}\big{/}\Delta_{\gamma}^{(t-1)}$ and keep only the first two terms in (9):
@@
4.2.2 KL Projection Stopping Criteria
Our assignment $\varepsilon_{\mathrm{d}}\propto\gamma^{-p}$ for some $p\geq 1$ in L4 of MDOT departs from the conventional wisdom of choosing $\varepsilon_{\mathrm{d}}\propto\gamma^{-1}$ (Altschuler et al., 2017; Dvurechensky et al., 2018; Lin et al., 2019). Here, we provide justification for this departure. Consider first the fixed-temperature problem (2.1) for simplicity. Building on the results of Cominetti & Martín (1994), Weed (2018) showed in his Prop. 4 and Thm. 5 that there is both a uniform bound $\langle P^{*}(\gamma)-P^{*},C\rangle\leq\log n/\gamma$ (slow rate), and a fast asymptotic rate $O(\exp(-\gamma K))$ which takes over for large enough $\gamma$, where the constant $K>0$ is problem-dependent. Taking these as a starting point, the following remark generalizes the third statement of Prop. 4.2.
@@ 4.2.2 KL Projection Stopping Criteria
Figure 3: Ablation of stopping criterion parameter $p$ (Sec. 4.2.2) for the SK algorithm (left) and MDOT with Sinkhorn and PNCG as KL projectors (right). The SK algorithm (left) is called by running MDOT (Alg. 1) with $\gamma_{\mathrm{i}}=\gamma_{\mathrm{f}}$, where higher precision is achieved by increasing $\gamma_{\mathrm{f}}$. Results show the median over 36 random problems from the upsampled MNIST dataset ($n=4096$) with the $L_{1}$ cost.
This result is consistent with the empirical findings of Jambulapati et al. (2019), who noted “The [tuned] Sinkhorn algorithm converged at rates much faster than the predicted $\varepsilon^{-2}$ rate on all experiments, outperforming all other methods, which we believe merits further investigation.” We believe Prop. 4.5 sheds some light on this phenomenon, and further present an ablation of $p$ in Fig. 3 for SK and MDOT algorithms.
For the SK algorithm, Fig. 3 (left) verifies the insight derived from Prop. 4.5. The choice $p=1$ is better at low precision, but the trend gradually shifts in favor of higher $p$ with (sufficiently) higher $\gamma_{\mathrm{f}}$. That is, for sufficiently low temperature $\gamma^{-1}$, it is advantageous to reduce the gradient norm error tolerance, from $H_{\min}/\gamma$ to $H_{\min}/\gamma^{p}$ for $p>1$. In contrast, MDOT is more robust to the $p$ parameter in the high precision regime (right). Moreover, the use of PNCG projections (Alg. 2) for KL projections in MDOT (L6 of Alg. 1) provides a speedup of 23×2-3$\times$2 - 3 × over Sinkhorn projections. PNCG is introduced and discussed next.
4.3 Preconditioned Non-linear Conjugate Gradients for KL Projections
SK converges more slowly at low temperatures (Kosowsky & Yuille, 1994). For a faster alternative, we develop Alg. 2 based on non-linear CG (NCG) methods (Fletcher & Reeves, 1964; Nocedal & Wright, 2006), which we now briefly review. Given an objective $g$, NCG takes descent directions ${{\bm{p}}^{(0)}=-\nabla g({\bm{z}}^{(0)})}$ and ${{\bm{p}}^{(k)}\leftarrow-\nabla g^{(k)}+\beta^{(k)}{\bm{p}}^{(k-1)}}$, and iterates ${{\bm{z}}^{(k+1)}\leftarrow{\bm{z}}^{(k)}+\alpha^{(k)}{\bm{p}}^{(k)}}$, where $\alpha^{(k)}$ is the step size. Optimal $\alpha^{(k)}$ has a closed-form solution for quadratics, but for general non-linear objectives, line search is necessary to find suitable step sizes $\alpha^{(k)}$. Many formulas for computing $\beta^{(k)}$ exist; for quadratic objectives, they are equivalent and guarantee convergence in at most $n^{\prime}$ iterations, where ${n^{\prime}\leq n}$ is the number of distinct eigenvalues of $\nabla^{2}g$. Further, the objective decreases faster if eigenvalues are tightly clustered (Stiefel, 1958; Kaniel, 1966; Nocedal & Wright, 2006). For example, the Hestenes-Stiefel formula sets (Nocedal & Wright, 2006):
$\begin{split}\tilde{{\bm{s}}}&=-\mathbf{D}\big{(}\mathbf{diag}(\nabla^{2}g)\big{)}^{-1}\nabla g=\left(\frac{{\bm{r}}}{{\bm{r}}(P)},\frac{{\bm{c}}}{{\bm{c}}(P)}\right)-\bm{1}_{2n}.\end{split}$ (14)
Algorithm 2 PNCG$({\bm{z}},\gamma,C,{\bm{r}},{\bm{c}},\varepsilon_{\mathrm{d}})$